OpenStudy (anonymous):

what about the inverse for the one to one function of.... f(x)= 2x-8

OpenStudy (amistre64):

switch out your y and x then solve for y x = 2y-8 x+8 ---- = y 2

OpenStudy (anonymous):

the answer to this one is [x+8 \over 2 \] =1/2x+4

OpenStudy (amistre64):

?

OpenStudy (anonymous):

x+8 over 2=1/2x+4....

OpenStudy (anonymous):

that's what my answer should be

OpenStudy (amistre64):

i see it; they simplified/modified it more...

OpenStudy (amistre64):

both are correct; just need to know how your program wants it formated

OpenStudy (anonymous):

im just studying for my final tomorrow

OpenStudy (amistre64):

which is correct; a $10 bill or ten $1 bills?

OpenStudy (anonymous):

they're the same. depends on the context.